# Towards Real-World Blind Face Restoration with Generative Diffusion Prior

### 25 Dec 2023 v1

@@ Abstract
Blind face restoration is an important task in computer vision and has gained significant attention due to its wide-range applications. In this work, we delve into the potential of leveraging the pretrained Stable Diffusion for blind face restoration. We propose BFRffusion which is thoughtfully designed to effectively extract features from low-quality face images and could restore realistic and faithful facial details with the generative prior of the pretrained Stable Diffusion. In addition, we build a privacy-preserving face dataset called PFHQ with balanced attributes like race, gender, and age. This dataset can serve as a viable alternative for training blind face restoration methods, effectively addressing privacy and bias concerns usually associated with the real face datasets. Through an extensive series of experiments, we demonstrate that our BFRffusion achieves state-of-the-art performance on both synthetic and real-world public testing datasets for blind face restoration and our PFHQ dataset is an available resource for training blind face restoration networks. The codes, pretrained models, and dataset are released at https://github.com/chenxx89/BFRffusion.
@@ I Introduction
Because of the unique structural and semantic information of face images, previous blind face restoration methods typically exploit face priors to restore face images, such as reference prior, geometric prior, and generative prior. Specifically, reference prior-based methods usually employ the facial structure [1, 2] or facial component dictionary [3, 4, 5] obtained from additional high-quality face images as the reference prior to guide the face restoration process. The unique geometric shapes and spatial distribution information of faces in the images are utilized to gradually recover high-quality face images in geometric prior-based methods. Geometric prior mainly include facial landmarks [6, 7], facial heatmaps [8], and facial parsing maps [9]. With the development of generative adversarial networks (GANs), researchers have started to leverage the generative prior for face image restoration. Generative prior typically includes GAN inversion [10] and pretrained facial GAN models [11, 12, 13] to provide richer and more diverse facial information.
@@
In this paper, we further explore the generative ability of the pretrained Stable Diffusion in the field of blind face restoration. The pretrained Stable Diffusion can provide rich and diverse priors including facial components and general object information, making it possible to generate realistic and faithful facial details. However, Stable Diffusion is a text-to-image generation model and is difficult to be applied to the restoration tasks directly. To address this, we propose BFRffusion with delicately designed architecture to leverage generative priors encapsulated in the pretrained Stable Diffusion for blind face restoration. Specifically, our BFRffusion consists of four modules: shallow degradation removal module (SDRM), multi-level feature extraction module (MFEM), trainable time-aware prompt module (TTPM), and pretrained denoising U-Net module (PDUM). SDRM removes the shallow degradation of the input and encodes it into latent space. MFEM adopts transformer blocks to extract multi-scale features added to the pretrained U-Net to guide the restoration process. TTPM generates the time-aware prompt which can provide semantic guidance for the restoration in different time steps. PDUM serves as the core denoising network of the diffusion model. Comprehensive experiments show that our BFRffusion achieves state-of-the-art performance on both synthetic and real-world datasets.
@@
Previous works [1, 9, 11, 4, 3, 5] mainly train the restoration networks on Flickr-Faces-HQ (FFHQ) [19] dataset. This dataset consists of 70K high-quality aligned face images collected from the Internet. However, real face datasets suffer from privacy issues and racial bias problems. What is more, collecting a high-quality dataset is a costly process. Recently, synthesis face datasets have received increasing attention in computer vision, and researchers [20, 21, 22, 23] have started utilizing these datasets for training neural networks. To address privacy and bias concerns of the real face datasets, we provide a synthetic face dataset called Privacy-preserving-Faces-HQ (PFHQ) with balanced race, gender, and age for training the restoration networks. Representative face images of our proposed PFHQ dataset are shown in Fig. I. To obtain the data set, we first employ ControlNet [24] which could add conditional information to the pretrained Stable Diffusion to generate a large number of face images. Then we select and classify the synthetic face images. Finally, we adopt a widely used degradation model to synthesize low-quality face images. Extensive experiments show that our PFHQ dataset achieves comparable performance in blind face restoration tasks compared to the widely used FFHQ dataset. We expect our PFHQ dataset can drive the development of face restoration and other face-related tasks in the field of computer vision.
@@
  • We leverage the generative priors encapsulated in the pretrained Stable Diffusion for blind face restoration. These priors contain abundant facial components and general object information, enabling us to restore realistic and faithful facial details.
  • We propose a blind face restoration method called BFRffusion with delicately designed architecture that could effectively extract multi-scale features from low-quality face images and guide the restoration process.
@@ II-A Diffusion models in Image Restoration
The latent diffusion model [16] proposes to implement diffusion-based generation in latent space to alleviate the training and sampling costs of the diffusion model. Inspired by this, StableSR [17] and DiffBIR [18] leverage the generative ability of the pretrained latent diffusion model to achieve realistic image restoration. DiffIR [15] exploits the latent-wise diffusion model to generate the compact image restoration priors, which guides the restoration network to achieve better performance.
@@ II-B Blind Face Restoration
Blind face restoration (BFR) aims to restore low-quality face images without any knowledge of degradation[35]. Compared with general images, the face image typically carries more structural and semantic information. Thus, most blind face restoration methods use facial priors to restore face images with clearer facial details.
Early attempts utilize reference prior [1, 36, 37] and geometric prior [38, 39, 8, 9] to guide the face restoration process. The geometric prior-based method usually extracted priors like facial landmarks [6, 7], facial heatmaps [8], and facial parsing maps [9] from low-quality face images. This also causes the restoration effect to be related to the degree of degradation. Reference prior-based methods [1, 36, 2, 40] guide the face restoration process by additional high-quality face images. Although this reduces the limitations of geometric priority, it may change the identity information of the restored face.
@@
Recently, the diffusion model has been proven to be more stable than GAN [25], and the generating images are more diverse. This has also received attention in the blind face restoration task. DR2 [44] diffuses input images to a noisy status where various types of degradation give way to Gaussian noise, and then captures semantic information through iterative denoising steps. DiffBIR [18] leverages the generative ability of pretrained Stable Diffusion as the prior and can achieve realistic image restoration.
@@ II-C Face Datasets
With the development of Artificial Intelligence Generated Content (AIGC), particularly the remarkable performance of diffusion models in the field of image generation in recent years, researchers [20, 21, 22, 23] have started utilizing synthetic face datasets to train neural networks. The Synthetic Faces High Quality (SFHQ) Dataset [50], consisting of approximately 425K synthetic images, was created by encoding the inspiration images into StyleGAN2 [42] latent space and then manipulating each image to make them appear photo-realistic. However, SFHQ is not paired and balanced. In order to address the ethical and imbalanced issues present in real face datasets and provide convenience to future blind face restoration research, we provide a privacy-preserving and paired face dataset with balanced race, gender, and age. Table I provides a summary of current face datasets to provide a comprehensive overview.
Figure 2: Overview of the architecture of the proposed BFRffusion. It comprises a shallow degradation removal module (SDRM), a multi-level feature extraction module (MFEM), a trainable time-aware prompt Module (TTPM), and a pretrained denoising U-Net module (PDUM).
@@ III-A1 Overall Architecture
As shown in Fig. 2, the proposed BFRffusion comprises four modules: shallow degradation removal module, multi-level feature extraction module, trainable time-aware prompt module, and pretrained denoising U-Net module. Specifically, given a degraded face image $x\in\mathbb{R}^{H\times W\times 3}$, the shallow degradation removal module, consisting of several convolutions, activation functions, and a ResBlock [51, 52], first encodes the input image $x$ into a latent representation $z\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times C}$ and extracts features $F_{1}$ from it. Then, the features $F_{1}$ are processed by the multi-level feature extraction module to extract multi-scale features that are appropriate for the different resolutions of Stable Diffusion. The multi-level feature extraction module is composed of several specially designed transformer blocks [53]. The trainable time-aware prompt module, constructed with a trainable parameter, a cross-attention block, and several multi-layer perceptron layers (MLP), generates the $Prompt$ that can guide the restoration process in different time steps. Finally, we add the output features $F^{n}$ from the multi-level feature extraction module to the pretrained denoising U-Net module to guide the denoising process and map the $Prompt$ from the trainable time-aware prompt module via a cross-attention layer to provide semantic guidance. A clear latent image is obtained from random Gaussian noise by gradual denoising and can be decoded to a clear image with the decoder of the pretrained VAE [54]. We provide a detailed explanation of the four modules in our BFRffusion in the subsequent sections.
@@ III-A2 Shallow Degradation Removal Module
III-A3 Multi-level Feature Extraction Module
Stable Diffusion is a typical U-Net architectural model that operates at four different resolutions: $64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$. To extract clear latent features from $F_{1}$ and align with the resolutions of Stable Diffusion, we propose a transformer-based U-Net called multi-level feature extraction module (MFEM). Our MFEM comprises several novel transformer blocks that consider both latent feature information and time conditions. We apply adaptive normalization operations [55, 56, 57] in the proposed transformer block to embed the time conditions. Specifically, we generate six affine transformation parameters $\alpha_{1},\beta_{1},\gamma_{1},\alpha_{2},\beta_{2},\gamma_{2}$ from the time embedding $emb$ in Eq. (2) using several MLP layers. After that, we apply Spatial Feature Transform (SFT) [58] to modulate the input image features that have undergone LayerNorm layer processing. It is formulated as follows:
@@ III-A4 Trainable Time-aware Prompt Module
Stable Diffusion 2.1-base is a large text-to-image diffusion model, which encodes the text information into vectors using a fixed, pretrained OpenCLIP-ViT/H text encoder with $354$M parameters. Previous works [17, 18] that utilize Stable Diffusion as the prior in Low-level Vision tasks usually set the prompt as null to generate the fixed latent vectors. However, there are two disadvantages by doing so: 1) The fixed CLIP [60] text encoder is so large that it requires a lot of computational resources during both the training and inference stages. 2) The operation is not beneficial to the tasks at all. To address these two problems, we propose the trainable time-aware prompt module (TTPM) to generate latent prompts that can guide the restoration process in different time steps. Specifically, our prompt component is a trainable parameter P $\in\mathbb{R}^{77\times 1024}$, which matches the size of the output of the CLIP text encoder. In order to generate more effective prompts in different time steps, we further incorporate the time embedding into the prompt via a cross-attention layer. The overall process of TTPM is defined as:
@@ III-B3 Stage III: Select and Classify the Face Images
Utilizing OpenCV, we measure the sharpness of synthetic face images via Laplacian variance and choose images with a sharpness surpassing 150 for subsequent classification. Fairface [49] is an artificially annotated dataset with balanced race, gender, and age. The prediction model trained on the Fairface dataset could achieve better classification accuracy. Thus, we employ this prediction model to classify the race, gender, and age of the selected images. We obtain four race groups: Asian, Black, White, and Other (including Indian, Latino, etc), two gender groups: male and female, and six age groups: 0-9, 10-19, 20-29, 30-39, 40-49 and 50+ years old. We randomly choose 1,250 face images from each distinct race, gender, and age group.
@@ III-B4 Stage IV: Synthesize Pairs of Images
Through the described procedures, we have obtained a dataset of 60K high-quality face images that are privacy-preserving and balanced. In order to fulfill the training requirements of current blind face restoration methods, we synthesize low-quality images from the corresponding high-quality images. We employ the degeneration strategies employed in most existing blind face restoration methods [63, 64, 1, 9, 11, 4, 3, 5, 18]. To be specific, we employ the following formula to synthesize the low-quality images from the high-quality images:
@@
In summary, by conducting the above four steps, we obtain a synthetic and balanced dataset comprising 60K pairs of face images that can be used for training blind face restoration methods.

### 18 Mar 2024 v2

@@ Abstract
Blind face restoration is an important task in computer vision and has gained significant attention due to its wide-range applications. Previous works mainly exploit facial priors to restore face images and have demonstrated high-quality results. However, generating faithful facial details remains a challenging problem due to the limited prior knowledge obtained from finite data. In this work, we delve into the potential of leveraging the pretrained Stable Diffusion for blind face restoration. We propose BFRffusion which is thoughtfully designed to effectively extract features from low-quality face images and could restore realistic and faithful facial details with the generative prior of the pretrained Stable Diffusion. In addition, we build a privacy-preserving face dataset called PFHQ with balanced attributes like race, gender, and age. This dataset can serve as a viable alternative for training blind face restoration networks, effectively addressing privacy and bias concerns usually associated with the real face datasets. Through an extensive series of experiments, we demonstrate that our BFRffusion achieves state-of-the-art performance on both synthetic and real-world public testing datasets for blind face restoration and our PFHQ dataset is an available resource for training blind face restoration networks. The codes, pretrained models, and dataset are released at https://github.com/chenxx89/BFRffusion.
@@ I Introduction
Because of the unique structural and semantic information of face images, previous blind face restoration methods typically exploit face priors to restore face images, such as reference prior [1, 2, 3], geometric prior [4, 5, 6, 7], and generative prior [8, 9, 10, 11]. Specifically, reference prior-based methods usually employ the facial structure [1, 2] or facial component dictionary [3] obtained from additional high-quality face images as the reference prior to guide the face restoration process. The unique geometric shapes and spatial distribution information of faces in the images are utilized to gradually recover high-quality face images in geometric prior-based methods. Geometric prior mainly include facial landmarks [4, 5], facial heatmaps [6], and facial parsing maps [7]. With the development of generative adversarial networks (GANs), researchers have started to leverage the generative prior for face image restoration. Generative prior typically includes GAN inversion [8] and pretrained facial GAN models [9, 10, 11] to provide richer and more diverse facial information. However, the prior knowledge derived solely from limited data may be not enough for restoration purposes.
@@
In this paper, we further explore the generative ability of the pretrained Stable Diffusion in the field of blind face restoration. Compared with the GAN prior used in previous generative prior-based methods [9, 10, 11], the pretrained Stable Diffusion can provide richer and more diverse priors including facial components and general object information, making it possible to generate realistic and faithful facial details. However, Stable Diffusion is a text-to-image generation model and can not be applied to restoration tasks directly without any modification. Although StableSR [15] and DiffBIR [16] leveraging the priors use existing U-net architecture of Stable Diffusion to extract image features, such design limits the restoration performance and efficiency. To address this, we propose BFRffusion with delicately designed architecture to leverage generative priors encapsulated in the pretrained Stable Diffusion for blind face restoration. Specifically, our BFRffusion consists of four modules: shallow degradation removal module (SDRM), multi-scale feature extraction module (MFEM), trainable time-aware prompt module (TTPM), and pretrained denoising U-Net module (PDUM). SDRM removes the shallow degradation of the input and encodes it into latent space. MFEM adopts transformer blocks to extract multi-scale features added to the pretrained U-Net to guide the restoration process. TTPM generates the time-aware prompt which can provide semantic guidance for the restoration in different time steps. PDUM serves as the core denoising network of the diffusion model. Compared with previous methods [15, 16] based on Stable Diffusion, more effective transformer blocks and time-aware prompts are applied in our BFRffusion. What is more, we employ a distinctive training strategy for the PDUM. Comprehensive experiments show that our BFRffusion achieves state-of-the-art performance on both synthetic and real-world datasets.
@@
Previous works [1, 7, 9, 17, 3, 18] mainly train the restoration networks on Flickr-Faces-HQ (FFHQ) [19] dataset. This dataset consists of 70K high-quality aligned face images collected from the Internet. However, real face datasets suffer from privacy issues and racial bias problems. What is more, collecting a high-quality dataset is a costly process. Recently, synthesis face datasets have received increasing attention in computer vision, and researchers [20, 21, 22, 23, 24] have started utilizing these datasets for training neural networks. To address privacy and bias concerns of the real face datasets, we provide a synthetic face dataset called Privacy-preserving-Faces-HQ (PFHQ) with balanced race, gender, and age for training the restoration networks. Representative face images of our proposed PFHQ dataset are shown in Fig. 1. To obtain the dataset, we first employ ControlNet [25] which could add conditional information to the pretrained Stable Diffusion to generate a large number of face images. Then we select and classify the synthetic face images. Finally, we adopt a widely used degradation model to synthesize low-quality face images. Meanwhile, we build PFHQ-Test using the same operations to evaluate the performance of restoration networks on balanced data. Extensive experiments show that our PFHQ dataset achieves comparable performance in blind face restoration tasks compared to the widely used FFHQ dataset. We expect our PFHQ dataset can drive the development of face restoration and other face-related tasks in the field of computer vision.
@@
  • We leverage the generative prior encapsulated in the pretrained Stable Diffusion for blind face restoration. The prior which contains abundant facial components and general object information is one of the prerequisites enabling us to restore realistic and faithful facial details.
  • We propose a blind face restoration method called BFRffusion with delicately designed architecture that could effectively extract multi-scale features from low-quality face images and sufficiently leverage the diffusion prior.
@@ II-A Diffusion models in Image Restoration
The latent diffusion model [14] proposes to implement diffusion-based generation in latent space to alleviate the training and sampling costs of the diffusion model. DiffIR [13] exploits the latent-wise diffusion model to generate the compact image restoration priors, which guides the restoration network to achieve better performance. StableSR [15] and DiffBIR [16] achieve realistic image restoration leveraging the generative ability of the pretrained Stable Diffusion which is based on the latent diffusion model. Our BFRffusion also leverages Stable Diffusion, but we apply delicate design and a distinctive training strategy.
@@ II-B Blind Face Restoration
Face restoration is a specific type of image restoration and the restored faces can be used for face recognition [36, 37, 38], face detection [39, 40] and etc. Blind face restoration (BFR) aims to restore low-quality face images without any knowledge of degradation [41]. Compared with general images, the face image typically carries more structural and semantic information. Thus, most blind face restoration methods use facial priors to restore face images with clearer facial details.
Early attempts utilize reference prior [1, 2, 42, 43] and geometric prior [4, 5, 44, 45, 6, 7] to guide the face restoration process. The geometric prior-based method usually extracted priors like facial landmarks [4, 5], facial heatmaps [6], and facial parsing maps [7] from low-quality face images. This also causes the restoration effect to be related to the degree of degradation. Reference prior-based methods [1, 42, 2, 46] guide the face restoration process by additional high-quality face images. Although this reduces the limitations of geometric priority, it may change the identity information of the restored face.
@@
Recently, the diffusion model has been proven to be more stable than GAN [26], and the generating images are more diverse. This has also received attention in the blind face restoration task. DR2 [50] diffuses input images to a noisy status where various types of degradation give way to Gaussian noise, and then captures semantic information through iterative denoising steps. DiffBIR [16] and our BFRffusion both leverage the pretrained Stable Diffusion as the generative prior which can provide more prior knowledge than other existing methods. Different from DiffBIR, we use delicately designed transformer architecture to extract multi-scale features and design a time-aware prompt module to replace the original CLIP used in DiffBIR to reduce the computational complexity.
@@ II-C Face Datasets
Figure 2: Overview of the architecture of BFRffusion which consists of four modules. The shallow degradation removal module (SDRM) and the multi-scale feature extraction module (MFEM) remove shallow degradation and extract multi-scale features from low-quality face images. The pretrained denoising U-Net module (PDUM) utilizes multi-scale features and prompts from the trainable time-aware prompt module (TTPM) as conditions to predict the next step of noise based on the input noise. After multiple denoising steps, high-quality latent features are obtained, which are subsequently transformed into high-quality face images by the pretrained decoder. The MFEM is composed of several transformer blocks, whose structure is illustrated below the dashed line.
With the development of Artificial Intelligence Generated Content (AIGC), particularly the remarkable performance of diffusion models in the field of image generation in recent years, researchers [20, 21, 22, 23, 24] have started utilizing synthetic face datasets to train neural networks. The Synthetic Faces High Quality (SFHQ) Dataset [56], consisting of approximately 425K synthetic images, was created by encoding the inspiration images into StyleGAN2 [48] latent space and then manipulating each image to make them appear photo-realistic. However, SFHQ is not paired and balanced. In order to address the ethical and imbalanced issues present in real face datasets and provide convenience to future blind face restoration research, we provide a privacy-preserving and paired face dataset with balanced race, gender, and age. Table I provides a summary of current face datasets to provide a comprehensive overview.
@@ III-A1 Overall Architecture
As shown in Fig. 2, the proposed BFRffusion comprises four modules: shallow degradation removal module (SDRM), multi-scale feature extraction module (MFEM), trainable time-aware prompt module (TTPM, and pretrained denoising U-Net module (PDUM). Specifically, given a degraded face image $x\in\mathbb{R}^{H\times W\times 3}$, the SDRM consisting of several convolutions, activation functions, and a ResBlock [57, 58], first encodes the input image $x$ into a latent representation $z\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times C}$ and extracts features $F_{1}$ from it. Then, the features $F_{1}$ are processed by the MFEM to extract multi-scale features that are appropriate for the different resolutions of Stable Diffusion. The MFEM is composed of several specially designed transformer blocks [59]. The TTPM constructed with a trainable parameter, a cross-attention block, and several multi-layer perceptron layers (MLP), generates the $Prompt$ that can guide the restoration process in different time steps. Finally, we add the output features $F^{n}$ from the MFEM to the PDUM to guide the denoising process and map the $Prompt$ from the TTPM via a cross-attention layer to provide semantic guidance. A clear latent image is obtained from random Gaussian noise by gradual denoising and can be decoded to a clear image with the decoder of the pretrained VAE [60]. We provide a detailed explanation of the four modules in our BFRffusion in the subsequent sections.
@@ III-A2 Shallow Degradation Removal Module
III-A3 Multi-scale Feature Extraction Module
Stable Diffusion is a typical U-Net architectural model that operates at four different resolutions: $64\times 64$, $32\times 32$, $16\times 16$, and $8\times 8$. To extract clear latent features from $F_{1}$ and align with the resolutions of Stable Diffusion, we propose a transformer-based U-Net called multi-scale feature extraction module (MFEM). Our MFEM comprises several novel transformer blocks that consider both latent feature information and time conditions. We apply adaptive normalization operations [61, 62, 63] in the proposed transformer block to embed the time conditions. Specifically, we generate six affine transformation parameters $\alpha_{1},\beta_{1},\gamma_{1},\alpha_{2},\beta_{2},\gamma_{2}$ from the time embedding $emb$ in Eq. (2) using several MLP layers. After that, we apply Spatial Feature Transform (SFT) [64] to modulate the input image features that have undergone LayerNorm layer processing. It is formulated as follows:
@@ III-A3 Multi-scale Feature Extraction Module
Figure 3: Visualization of feature maps learned by our multi-scale feature extraction module (MFEM) in different timesteps and resolutions. The first row demonstrates the capability of our MFEM to extract accurate features at any timesteps. The second row shows the multi-scale features extracted by our MFEM at various resolutions.
Fig. 3 shows the visualization of $F^{n}$ learned by our MFEM in different timesteps and resolutions. Thanks to the delicately designed transformer blocks, our MFEM can extract accurate features at any timesteps. Besides, we find that the feature maps around 200 timesteps are clearer, and such phenomena have similarly occurred in previous works [15]. The second row shows the multi-scale features that match the different resolutions of Stable Diffusion blocks. The proposed MFEM constitutes one of the essential prerequisites enabling our method to achieve state-of-the-art performance.
@@ III-A4 Trainable Time-aware Prompt Module
Stable Diffusion 2.1-base is a large text-to-image diffusion model, which encodes the text information into vectors using a fixed, pretrained OpenCLIP-ViT/H text encoder with $354$M parameters. Previous works [15, 16] that utilize Stable Diffusion as the prior in Low-level Vision tasks usually set the prompt as null to generate the fixed latent vectors. However, there are two disadvantages by doing so: 1) The fixed CLIP [66] text encoder is so large that it requires a lot of computational resources during both the training and inference stages. 2) The operation is not beneficial to the tasks at all.
To address these two problems, we propose the trainable time-aware prompt module (TTPM) to generate latent prompts that can guide the restoration process in different time steps. Specifically, our prompt component is a trainable parameter P $\in\mathbb{R}^{77\times 1024}$, which matches the size of the output of the CLIP text encoder. In order to generate more effective prompts in different time steps, we further incorporate the time embedding into the prompt via a cross-attention layer. The overall process of TTPM is defined as:
@@
TABLE II: The computational complexity between our BFRffusion and other diffusion-based methods
Method Parameters MACs Time(s/image)
StableSR [15] 1.20B 1498.52G 1.84
DiffBIR [16] 1.51B 1632.51G 2.97
BFRffusion(ours) 1.23B 986.38G 1.45
The computational complexity between our BFRffusion and other diffusion-based methods is shown in Table II. We utilize the profile function from the thop package to compute the parameter count and Multiply-Accumulate Operations (MACs). The inference time is conducted on 1 NVIDIA A100 GPU. Thanks to our TTPM, our method significantly reduces computational complexity and holds an advantage in inference time.
@@ III-B3 Stage III: Select and Classify the Face Images
Utilizing OpenCV, we measure the sharpness of synthetic face images via Laplacian variance and choose images with a sharpness surpassing 150 for subsequent classification. Fairface [55] is an artificially annotated dataset with balanced race, gender, and age. The prediction model trained on the Fairface dataset could achieve better classification accuracy. Thus, we employ this prediction model to classify the race, gender, and age of the selected images. We obtain four race groups: Asian, Black, White, and Other (including Indian, Latino, etc), two gender groups: male and female, and six age groups: 0-9, 10-19, 20-29, 30-39, 40-49 and 50+ years old. We randomly choose 1,250 and 10 face images to build the training and testing dataset from each distinct race, gender, and age group.
@@ III-B4 Stage IV: Synthesize Pairs of Images
Through the described procedures, we have obtained 60K and 480 high-quality face images that are both privacy-preserving and balanced. In order to fulfill the training and testing requirements of current blind face restoration methods, we synthesize low-quality images from the corresponding high-quality images. We employ the degeneration strategies employed in most existing blind face restoration methods [69, 70, 1, 7, 9, 17, 3, 18, 16]. To be specific, we employ the following formula to synthesize the low-quality images from the high-quality images:
@@
In summary, by conducting the above four steps, we obtain a synthetic and balanced training dataset comprising 60K pairs of face images that can be used for training blind face restoration methods. We use the other 480 pairs of face images to build the PFHQ-Test to evaluate the performance of restoration networks on balanced data.